# Seeing through Light and Darkness: Sensor-Physics Grounded Deblurring HDR NeRF from Single-Exposure Images and Events

### 13 Mar 2026 v3

@@ C.1 Exposure time settings
C.2 Pose settings
We use the pose generation framework of E2NeRF [35] to obtain the poses corresponding to the blurry images for all the compared methods and our See-NeRF, which ensures the input poses are the same, so the reconstruction result is only affected by the network architecture.
@@ C.3 Evaluation of HDR results
For the HDR NVS task, we first use the “Enhancer” and “Compressor” options of Photomatrix Pro [16] for our synthetic and real data, respectively, to tonemap the reconstructed and ground truth HDR images into the LDR domain. Then we use PSNR, SSIM, and LPIPS [52] to quantitatively and qualitatively show the results in our main paper and supplementary materials, which are the same as [4, 19].
@@ D Additional evaluations
In this experiment, we train HDR-NeRF [19] with different input exposure ranges to evaluate the HDR NVS performance boundaries of multi-exposure image-based HDR-NeRF and our single-exposure ERGB-based See-NeRF.
@@
Quantitative results: HDR-NeRFref in Tab. 7 demonstrates that multi-exposure image-based HDR-NeRF can only fully reconstruct sharp HDR scenes when the input LDR images are sharp and cover exposure times at all three scales: low (), medium (), and high (). When the input long-exposure images are motion-blurred, the results of HDR-NeRFref-blur will be significantly affected by the motion blur and degraded. In contrast, our See-NeRF achieves the best HDR NVS results with blurry single-exposure LDR images and corresponding events
Qualitative results: HDR-NeRF in Fig. 9 illustrates that with images at single-exposure time as input, the estimated CRF curve deviates completely from GT, and the HDR NVS results are also completely distorted. When adding the images with longer exposure time for training, HDR-NeRF can accurately recover the CRF curve of the low-light area, while there is still a deviation in the CRF curve of the highlight area, and the HDR NVS results will show a color cast in the highlight area. When further adding the image with shorter exposure time , the CRF curve of the highlight area is further corrected and close to the GT CRF curve, as shown in the results of HDR-NeRFref. However, in the case of blurry input, the estimated CRF curve of HDR-NeRFref-blur will be distorted, and its reconstruction results will be blurred. In comparison, our See-NeRF leverages the spatial dynamic information of the low-light and highlight areas contained in the event to correct the CRF curves in these areas, achieving both HDR and deblurring NVS with a single-exposure image and corresponding events as input.
Therefore, See-NeRF effectively leverages the spatial and temporal differential information in the low-light, highlight, and blurry regions that are contained in the events to infer the actual brightness of the scene and eliminates the effects of the blurred input, respectively, achieving the best deblurring HDR NVS results with single-exposure input and events, as deduced in the Sec. 3 of the main paper.
@@
D.2 Parameters and training time evaluation
Parameter : in Eq.(6) represents the discretization of Eq.(1) with discrete time points. Therefore, a larger will lead to better simulation of image formation and generate better results, but at the same time require more iterations and increase the training time as shown in Tab. 8. For the HDR NVS task, when , See-NeRF gets the best results. For the deblurring NVS task, when , the improvement is marginal. So we set for the HDR and novel exposure NVS experiments on our proposed datasets and for the deblurring NVS experiments for the Real-World-Challenge dataset.
Training time: Compared to HDR-NeRF (2.1h), See-NeRF (2.5h) increases computation and training time by 19.05% under the same conditions (same NeRF-related parameters, number of iterations, and batch size; on a single RTX 3090 GPU; on our real dataset). The increased computation is mainly due to the introduction of events and calculations related to event mapping fields. However, since no events are involved during test time, the computation and rendering speed are close to HDR-NeRF (At 346*260 resolution, rendering each frame takes 2.5 seconds).
Parameter : As shown in Fig. 10, we evaluate the impact of in real-world “desktop” and “lab” scenes for the HDR and deblurring NVS tasks, respectively. When , See-NeRF can achieve the best overall results on both tasks. Note that represents train See-NeRF without event loss supervision.
@@ D.3 Sensor-Physics grounded deblurring effect
Our See-NeRF is significantly better than previous ERGB-based deblurring NVS methods E2NeRF [35], E3NeRF [34], E2GS [10], and Ev-DeblurNeRF [5] as in Tab. 16. This is because these methods all perform the motion blur and events synthesis directly on the nonlinear LDR images rendered by their NeRF network, as shown in Fig. 11. In contrast, our See-NeRF performs the motion blur and events synthesis first on the linear raw HDR values of the scene radiance rendered by the NeRF network, and then tonemaps the blurry raw HDR image to the LDR domain, aligning with the physical imaging process of sensors in the real world. On the other hand, our event mapping field eliminates the effects of event delay and further improves the performance. Therefore, See-NeRF achieves SOTA results in the ERGB-based deblurring NeRF task.
| Imaging Model | Event | Real Data HDR | Real Data Deblurring | ||||||
| Method | Integration | CRF | Non-Idealities | PSNR↑ | SSIM↑ | LPIPS↓ | PSNR↑ | SSIM↑ | LPIPS↓ |
| EvHDR-GS | N/A | N/A | 18.94 | .7124 | .2510 | 28.69 | .9174 | .2674 | |
| See-NeRF | Eq. (6) | MLP-Based RGB CRF | Sec. 4.3 | 26.49 | .8953 | .1638 | 32.70 | .9564 | .1574 |
| See-NeRF# | N/A | Sec. 4.3 | 19.75 | .7239 | .2696 | 28.35 | .9242 | .2657 | |
D.4 Discussion with EvHDR-GS
@@
A recent work, EvHDR-GS [7], uses 3D gaussian splatting as a backbone for the ERGB-based HDR video reconstruction task. Though it uses a similar design that RGB CRF is applied after volumetric rendering, it does not consider the image integration process and simply models the RGB CRF with one exponential function during the 2D mapping, while See-NeRF address this two aspects with Eq. (6) in the main paper and a three-independent MLP-Based RGB CRF model as shown in the Fig. 3 of the main paper. Besides, we model the event non-idealities with our event mapping fields. Therefore, EvHDR-GS underperforms See-NeRF on both deblurring and HDR tasks, and applying its imaging model to See-NeRF also degrades the performance significantly, as shown in Tab. 9.
D.5 Pose estimation
We use the event-enhanced pose estimation strategy of E2NeRF for our See-NeRF. Recent image-based deblurring NVS works provide different camera pose estimation strategies [43, 54, 24] and pose change estimation strategies for dynamic scenes [3, 25]. Integrating the event supervision into these advanced pose estimation strategies, as in EBAD-NeRF [36], would be a promising future research direction for the ERGB-based HDR deblurring NVS task.
E Detailed results
E.1 Detailed quantitative results
Our synthetic dataset: Tab. 10 and Tab. 11 illustrate the detailed quantitative novel exposure NVS results on each scene of our synthetic dataset. See-NeRF achieves the best overall performance. In certain scenes, the novel exposure results of the compared RGB-based HDR NVS methods are better than those of See-NeRF because it uses sharp images as input. Besides, the detailed quantitative results in Tab. 12 and Tab. 13 illustrate that their performance drops significantly on the HDR NVS task, indicating that they are overfit to the novel exposure task, thus learning a wrong scene HDR representation. In contrast, See-NeRF achieves the best HDR NVS results for all three metrics.
Our real dataset: Tab. 14 and Tab. 15 illustrate the detailed quantitative novel exposure and HDR NVS results on each scene of our real dataset. Some results of the compared RGB-based HDR NVS methods are better than those of See-NeRF on LPIPS due to the sharp images input, while our See-NeRF still achieves the best overall performance. For the HDR experiment, See-NeRF is even better than HDR-NeRFref in “Bear”, “Computerhard”, and “Desktop” scenes, resulting in the average results in Tab. 1 of the main paper outperforms HDR-NeRFref. This is likely attributable to the low imaging quality of the DAVIS 346 RGB sensor, which is not suitable for image-based methods.
E.2 Detailed qualitative results
Our synthetic dataset: Fig. 12 shows the detailed qualitative HDR NVS results on our synthetic dataset, and the results of See-NeRF are closest to GT across all scenes.
Our real dataset Fig. 13 shows the detailed qualitative HDR NVS results on our real dataset, and the results of See-NeRF are closest to GT across all scenes.
Real-World-Challenge dataset Fig. 14 shows the detailed qualitative deblurring NVS results on the Real-World-Challenge dataset. See-NeRF shows impressive deblurring performance compared to other methods.
E.3 Supplementary video results
We also provide supplementary video rendering results to show the impressive HDR and Deblurring NeRF novel view synthesis results of See-NeRF and robustness to noise events of our framework. We highly recommend that the reviewers refer to the attached video for details.

### 02 Apr 2026 v4

@@ C.1 Exposure time settings
C.2 Pose estimation
We use the pose generation framework of E2NeRF [37] to obtain the poses corresponding to the blurry images for all the compared methods and our See-NeRF, which ensures the input poses are the same, so the reconstruction result is only affected by the network architecture. Recent RGB-based deblurring NVS works provide different camera pose estimation strategies [47, 58, 24] and dynamic scene pose estimation strategies [3, 25]. Integrating the event supervision into these methods would be a promising future research direction for the ERGB-based HDR deblurring NVS task, as demonstrated in EBAD-NeRF [38].
@@ C.3 Evaluation of HDR results
For the HDR NVS task, we first use the “Enhancer” and “Compressor” options of Photomatrix Pro [16] to tone map the reconstructed and ground truth HDR images into the LDR domain for the synthetic and real data, respectively. Then, we use the tone-mapped images to quantitatively and qualitatively evaluate our results in our main paper and supplementary materials, as in [4, 19].
@@ D Additional evaluations
In this experiment, we train HDR-NeRF [19] with different input exposure ranges to evaluate the HDR NVS performance boundaries of multi-exposure RGB-based HDR-NeRF and our single-exposure ERGB-based See-NeRF.
@@
Quantitative results: HDR-NeRFref in Tab. 7 demonstrates that multi-exposure RGB-based HDR-NeRF only surpasses our single-exposure ERGB-Based See-NeRF in overall performance when the input LDR images are sharp and cover exposure times at all three scales: short (), medium (), and long (). When the input images are motion-blurred, the results of HDR-NeRFref-blur will be significantly affected by the motion blur and degraded. In contrast, our See-NeRF achieves the best HDR NVS results with blurry single-exposure LDR images and corresponding events.
Qualitative results: HDR-NeRF in Fig. 9 illustrates that with images with single-exposure time as input, the estimated CRF curve deviates completely from GT, and the HDR NVS results are also completely distorted. When adding the images with longer exposure time for training, HDR-NeRF can accurately recover the CRF curve of the low-light area. At the same time, there is still a deviation in the CRF curve of the highlight area, and the HDR NVS results will show a color cast in the highlight area. When further adding the images with shorter exposure time , the CRF curve of the highlight area is further corrected and closer to the GT CRF curve, as shown in the results of HDR-NeRFref. However, in the case of blurry input, the estimated CRF curves of HDR-NeRFref-blur are distorted, and the reconstruction results are blurred. In comparison, See-NeRF achieves accurate CRF curve estimation and produces sharp, high-quality HDR NVS results. This demonstrates that See-NeRF effectively leverages spatial and temporal differential information of events in low-light, highlight, and blurry regions to infer the actual scene brightness to enhance CRF estimation and eliminate the effects of blurred input, which is consistent with the theoretical proof in Sec. 3.
@@
| Our Real Data (HDR) | Real-World-Challenge Data | |||||||
| Event CRF Model | RGB CRF Model | PSNR↑ | SSIM↑ | LPIPS↓ | PSNR↑ | SSIM↑ | LPIPS↓ | |
| See-NeRFPow | ours | Power Function CRF [7] | 19.75 | .7239 | .2696 | 28.35 | .9242 | .2657 |
| See-NeRF | ours | MLP-Based CRF (ours) | 26.49 | .8953 | .1638 | 32.70 | .9564 | .1574 |
| See-NeRFv2e | Explicitly Parameterized 2nd-Order LPF [17] | ours | 22.10 | .8644 | .1938 | 30.51 | .9441 | .1748 |
| See-NeRFDeblur e | Explicitly Parameterized 4th-Order LPF [31] | ours | 24.28 | 8859 | .1847 | 31.84 | .9521 | .1664 |
| See-NeRF | Implicitly MLP-Optimized 2nd-Order LPF (ours) | ours | 26.49 | .8953 | .1638 | 32.70 | .9564 | .1574 |
D.2 Parameters and efficiency
Parameter : in Eq. (6) represents the discretization of Eq. (1) with discrete time points. Therefore, a larger leads to better simulation of the imaging process and generates better results, but at the same time requires more iterations and increases the training time as shown in Tab. 8. For the HDR NVS task, when , See-NeRF gets the best results. For the deblurring NVS task, when , the improvement is marginal. So we set for the HDR and novel exposure NVS experiments on our proposed real-world datasets and for the deblurring NVS experiments for the Real-World-Challenge dataset.
Training and rendering time: Compared to HDR-NeRF (2.1h), See-NeRF (2.5h) increases computation and training time by 19.05% under the same conditions (same NeRF-related parameters, number of iterations, and batch size; on a single RTX 3090 GPU; on our real dataset). The increased computation is mainly due to the event-related calculations. However, the rendering speed of See-NeRF (0.37 fps) is slightly faster than HDR-NeRF (0.35 fps) under a resolution of 346*260. This is mainly because no events are involved during test time, and our 2D pixel-wise tone mapping is more efficient than the 3D points tone mapping of HDR-NeRF.
Parameter : As in Fig. 10, we evaluate the impact of on the real-world “desktop” and “lab” scenes for the HDR and deblurring NVS tasks, respectively. When , See-NeRF can achieve the best overall results on both tasks.
@@ D.3 Sensor-Physics grounded deblurring effect
Our See-NeRF significantly surpasses previous ERGB-based deblurring NVS methods E2NeRF [37], E3NeRF [36], E2GS [10], and Ev-DeblurNeRF [5] as shown in Tab. 16. This is because these methods all perform the motion blur and events synthesis directly on the nonlinear LDR images rendered by their 3D representation network, as shown in Fig. 11. In contrast, our See-NeRF performs the motion blur and events synthesis on the linear raw HDR values of the scene radiance rendered by the NeRF network, and then tone maps the blurry raw HDR image to the LDR domain, aligning with the physical imaging process of sensors in the real world. On the other hand, our event CRF model eliminates the effects of event delay and further improves the performance. Therefore, See-NeRF achieves SOTA results in the ERGB-based deblurring NeRF task.
D.4 Ablation on previous related CRF models
We conduct an ablation study on previous related CRF models on our real-world dataset and the Real-World-Challenge dataset, respectively, as shown in Tab. 9
@@ D.4 Ablation on previous related CRF models
RGB CRF model of EvHDR-GS [7]: Recently, EvHDR-GS uses 3DGS [23] as a backbone for the ERGB-based HDR video reconstruction task. Though it uses a similar design that RGB CRF is applied after volumetric rendering, it does not consider the image integration process and simply models the RGB CRF with one power function for the three color channels during the 2D tone mapping, while See-NeRF address these two aspects with Eq. (6) in the main paper and a three-independent MLP-Based RGB CRF model, as shown in the Fig. 3 of the main paper. Therefore, applying its imaging model to See-NeRF degrades the performance significantly on both HDR and deblurring NVS tasks, as shown in the results of See-NeRFPow in Tab. 9.
Event CRF model of v2e [17] and Deblur e-NeRF [31]: v2e and Deblur e-NeRF models intensity-dependent pixel bandwidth and latency with the explicitly parameterized 2nd-order and 4th-order low-pass filter, respectively. However, our See-NeRF incorporates scene/camera-related event delay factors into an MLP to calculate the delay coefficient for the 2nd-order low-pass filter. Replacing our event CRF model with the models of v2e and Deblur e-NeRF weakens the performance of See-NeRF on both deblurring and HDR tasks, as in the results of See-NeRFv2e and See-NeRFDeblur e in Tab. 9, proving the effectiveness of our event CRF model.
@@
E Supplementary detailed results
E.1 Detailed quantitative results
Our synthetic dataset: Tab. 10 and Tab. 11 illustrate the detailed quantitative novel exposure NVS results on each scene of our synthetic dataset. See-NeRF achieves the best overall performance. In certain scenes, the novel exposure results of the compared RGB-based HDR NVS methods are better than those of See-NeRF because it uses sharp images as input. Besides, the detailed quantitative results in Tab. 12 and Tab. 13 illustrate that their performance drops significantly on the HDR NVS task, indicating that they are overfitted to the LDR image supervision, thus learning a wrong scene HDR representation. In contrast, See-NeRF achieves the best HDR NVS results for all three metrics.
Our real dataset: Tab. 14 and Tab. 15 illustrate the detailed quantitative novel exposure and HDR NVS results on each scene of our real dataset. Some results of the compared RGB-based HDR NVS methods are better than those of See-NeRF on LPIPS due to the sharp images input, while our See-NeRF still achieves the best overall performance. For the HDR experiment, See-NeRF is even better than HDR-NeRFref in “Bear”, “Computerhard”, and “Desktop” scenes, resulting in its average results in Tab. 1 of the main paper outperforming HDR-NeRFref. This is likely attributable to the low imaging quality of the DAVIS 346 RGB sensor, which is not suitable for RGB-based methods.
E.2 Detailed qualitative results
Our synthetic dataset: Fig. 12 shows the detailed qualitative HDR NVS results on our synthetic dataset. The results of See-NeRF are closest to GT across all scenes. HDR-GS suffers from severe color shifts. HDR-NeRF fails to reconstruct the overexposed and underexposed areas. HDR-NeRF+ is limited by the HDREv-Net [52], which only achieves a brightening effect relative to the input image without fully restoring the scene color details. EvHDR-NeRF still suffers from slight color casts and motion blur.
Our real dataset Fig. 13 shows the detailed qualitative HDR NVS results on our real dataset, and the results of See-NeRF are closest to GT across all scenes. The compared methods show the same phenomenon as in the synthetic data.
Real-World-Challenge dataset Fig. 14 shows the detailed qualitative deblurring NVS results on the Real-World-Challenge dataset. See-NeRF shows impressive deblurring performance compared to other methods. Note that EvHDR-NeRF is significantly affected by the motion blur.
E.3 Supplementary video results
We also provide supplementary video rendering results to show the impressive HDR and deblurring novel view synthesis results of See-NeRF and the robustness to noise events of our framework. We highly recommend that the reviewers refer to the attached video for details on our project web page: https://icvteam.github.io/See-NeRF.html.